# Decoding Geometric Properties in Non-Random Data from First Information-Theoretic Principles

### v2

@@ 4.2 Toward a general theory for zero-knowledge one-way communication
In the general case (where the multidimensional spaces may be much more complex than the objects themselves), even in the ideal scenario in which we correctly infer the actual object sent by a completely unknown emitter, we are still prone to being wrong about the original dimensions’ configurations the emitter intended [35, 36]. (See also Section A.2 ) This poses a challenge to straightforward applications of the algorithmic information theory (AIT). In order to tackle such a challenge, we develop a method for multidimensional reconstruction of received low-dimensional signals in which one does not have access to knowledge about the encoding-decoding scheme or the original multidimensional space. Theorem 1 demonstrates the existence of sufficient conditions (or assumptions) that allow one to achieve an encoding/decoding-agnostic method for reconstructing the original multidimensional spaces. It gives rise to an agnostic method that does not need to assume that one a priori knows that the emitter agent’s multidimensional space resembles ours, is as complex as ours, or is isomorphic to the structures we currently deem fundamental in mathematics.
Theorem 1 (Theorem 11 in the Sup. Inf.).
@@
hold, where $\mathcal{P}$ is any algorithmic perturbation that transforms $y$ embedded into $\mathcal{S}$ into another $y^{\prime}$ embedded into $\mathcal{S}^{\prime}$. Let $\mathbf{P}^{\prime}\left[\cdot,\cdot\right]$ be the uniform probability measure over the space of all possible outcomes (in the form $\left(y^{\prime},\,\mathcal{S}^{\prime}\right)$) of algorithmic perturbations on $\left(y,\,\mathcal{S}\right)$ that satisfy Equation 3. Let $\mathbf{P}\left[\cdot,\cdot\right]$ be any arbitrarily chosen (computable) probability measure on the space of all possible (computably constructible) objects embedded into a multidimensional space. Then, the following strong asymptotic dominance
@@
holds with probability (given by the uniform probability measure $\mathbf{P}^{\prime}$) arbitrarily close to $1$ as $\left|y\right|\to\infty$ for $\left(y^{\prime},\mathcal{S}^{\prime}\right)$ distinct from $\left(y,\mathcal{S}\right)$ satisfying Equation 3.
@@
As further discussed in Section A.5, we believe the results in this article not only contribute with a general method for interpreting messages in zero-knowledge one-way communication channels, but also establish fundamental connections (and posit open problems) with zero-knowledge proofs and the algorithmic counterparts of the maximum likelihood method and the noisy-channel coding theorem.
[…]
Acknowledgements
@@
Appendix A Supplementary Information
A.1 Introduction to the formal concepts
In what follows, we specify the scientific context and mathematical conditions that enable our results and claims.
For the present purposes, we say an object has “meaning” when the object conveyed in the received signal is actually embedded into (or grounded on) the multidimensional space (i.e., the context) that the emitter agent originally intended. Each (finite and discrete) multidimensional space $\mathcal{S}$—no matter how complex it is—can be univocally determined by the number of dimensions, the encoding of the set of elements for each dimension, and the ordering in which each dimension appears. For example: the unidimensional space $\mathcal{S}_{1}$ that takes values from the natural numbers can be univocally determined by only informing the length $\left|\mathcal{S}_{1}\right|$ in $\mathbf{O}\left(\log\left(\left|\mathcal{S}_{1}\right|\right)\right)$ bits; the bidimensional space $\mathcal{S}_{2}$ by a pair $\left(\left|{\mathcal{S}_{2}}_{x}\right|,\left|{\mathcal{S}_{2}}_{y}\right|\right)$ in $\mathbf{O}\left(\log\left(\left|{\mathcal{S}_{2}}_{x}\right|\right)\right)+\mathbf{O}\left(\log\left(\left|{\mathcal{S}_{2}}_{y}\right|\right)\right)+\mathbf{O}\left(1\right)$ bits, where ${\mathcal{S}_{2}}_{x}$ is the first dimension and ${\mathcal{S}_{2}}_{y}$ is the second dimension; and so on. In this manner, one is always able to (uniquely) decode the received message into sufficient information for completely determining the multidimensional space configuration. One example of encoding scheme for an arbitrary multidimensional space $\mathcal{S}$ is by encoding it in the form of a companion tuple $\bm{\tau}$ as in [35, 36]. Finite and discrete multidimensional spaces with more intricate configurations other than whole numbers multiplication in the form of $m_{1}\times m_{2}\times\dots\times m_{k}$, where $m,k\in\mathbb{N}$, can for example be encoded by node-unaligned companion tuples $\bm{\tau}_{ua}$ as in [36, Definition 2]. Nevertheless, as demonstrated in Section 2.2, the theoretical framework of our method is agnostic to any arbitrarily chosen computational scheme to encode multidimensional spaces.
Each partition of a $n$-bit-length linear signal stream corresponds to a distinct configuration of dimensions’ lengths, and therefore to a distinct multidimensional space in which the total additive dimensions’ lengths remains upper bounded by the constant $n$. For example, a picture of 4$\times$4 pixels can be partitioned into other dimension-length configurations leading to different non-squared shapes, like 2$\times$8, 8$\times$2, or 1$\times$16. Once a row (or, alternatively, a column) is fixed in a bidimensional space, the remaining rows and columns are dependent on the initial row fixation. In other words, each bidimensional partition is a single-variable dependency. More formally, in the case of a (finite and discrete) bidimensional space $\mathcal{S}_{2}$, one may have different combinations of values $\left|{\mathcal{S}_{2}}_{x}\right|$ and $\left|{\mathcal{S}_{2}}_{y}\right|$ for each partition. However, for any partition one has it that both $1\leq\left|{\mathcal{S}_{2}}_{x}\right|\leq n$ and $\lfloor\frac{n}{\left|{\mathcal{S}_{2}}_{x}\right|}\rfloor=\left|{\mathcal{S}_{2}}_{y}\right|$ hold. Thus, because each partition is single-variable dependent on ${\mathcal{S}_{2}}_{x}$, for any $n$-bit-length linear signal stream one can encode each of its $2$D-partitions in $\mathbf{O}\left(\log\left(n\right)\right)$ bits. The same applies to $3$D-partitions (or to any partition from a finite number of dimensions), except for an extra number of bits upper bounded by a partition-independent constant $\mathbf{O}\left(1\right)$, which basically corresponds to the algorithmic complexity of the minimal program that calculates the other dimensions given the respective single variable as input.
As formalised in [26] any transformation or perturbation of a finite (encoded) object into another finite (encoded) object is equivalent to an algorithmic transformation that takes the former as input and outputs the latter. (See also Section A.3.2). Thus, a (algorithmic) perturbation with low algorithmic complexity is one of these transformations whose algorithmic information content of the original object is mostly preserved under such perturbations. More formally, a low-complexity perturbation occurs when $\mathbf{K}\left(y\middle|x\right)=\mathbf{O}\left(\log\left(\left|x\right|\right)\right)$, where $\mathbf{K}\left(\cdot\right)$ denotes the algorithmic complexity, $x$ is the original encoded object, $y$ is the encoded object that results from the perturbation on $x$, and $\left|x\right|$ is the size of the object $x$.
Swapping all 0’s and 1’s, or swapping all black and white pixels in an image are examples of low-complexity perturbations (in these two cases in particular, one has it that $\mathbf{K}\left(y\middle|x\right)=\mathbf{O}\left(1\right)$). Randomly inserting or deleting a finite number of elements (such as creating or destroying edges in a graph or flipping a finite number of bits in a string) are also examples of low-complexity perturbations [26, 37, 24]. As we saw in the above paragraphs, changing the partition is also a low-complexity perturbation of the multidimensional space into which the object $x$ is embedded. This is because any new partition can always be encoded in $\mathbf{O}\left(\log\left(\left|x\right|\right)\right)$ bits. Thus, notice that a perturbation can change either the object (e.g., by flipping bits) or its multidimensional space (e.g., by reconfiguring the dimensions’ lengths of a given partition to achieve another distinct partition) independently, or can change both at the same time. In any of these three options, the encoded form of the message resulting from the perturbation will be distinct from the encoded form of the original message. Thus, notice that both changing the partition into which the object is embedded and introducing noise into the object itself (e.g., by flipping bits) are particular examples of (algorithmic) perturbations [26, 37, 24].
The theoretical framework and results presented in Section 2.2 demonstrate the method employed to achieve the empirical findings in Section A.4 is sound in the infinite asymptotic limit when computational resources are unbounded, and it can be generalised to multidimensional spaces in addition to bi- or tridimensional ones. Our empirical results corroborate the theoretical predictions of the theorems in Section 2.2, which in turn evinces that AID is capable of harnessing (under feasible computational resources) some of the algorithmic-informational properties that the following theorems demonstrate in ideal and generalised scenarios. As shown in [24, 32], AID allows the investigation of causal analysis [28, 29] and solution of inverse problems [25] by taking advantage of a high convergence rate to the algorithmic probability brought along by BDM, probability values which remain stable across distinct models of computation, and most prominently for high algorithmically probable objects.
In Section A.2, we present the theoretical limitations in the general case, which in turn highlights the sufficient conditions that should be satisfied in order to avoid these limitations as we will investigate in the theory developed in Section 2.2.
@@ A.1 Introduction to the formal concepts
A.2 Avoiding algorithmic information distortions in arbitrarily complex multidimensional spaces
@@
In [35, 36] it is investigated how much of the irreducible information content of a multidimensional network and its isomorphic monoplex network (i.e., its isomorphic graph) is preserved during such an isomorphism transformation. In fact, isomorphisms are demonstrated to not preserve algorithmic information in the general case. The algorithmic information of a multidimensional network might be exponentially distorted with respect to the algorithmic information of its isomorphic (monoplex) network (see [36, Theorem 12]). We refer to an algorithmic information distortion when the algorithmic information content of an object (e.g., measured by the prefix algorithmic complexity of an encoded form of the object) in a certain context (e.g., the multidimensional space the object is embedded into) sufficiently differs from the algorithmic information content of the same object in a distinct context. In this regard, algorithmic complexity oscillations may arise from a series of distortions produced by a large enough collection of distinct contexts into which the same object is embedded.
@@
As shown in Section A.4, other examples are the oscillations measured by the plots presented in the figures, particularly those image plots for which one has the $x$-axis representing the distinct partitions we estimated the algorithmic complexity of the messages into, respectively. Notice that the downward spikes one can see in Section A.4 are examples of algorithmic complexity (negative) oscillations such that for the decoding problem we are interested in, they indicate the partition that corresponds to the correct multidimensional space the message intended.
@@
The results in [35, 36] demonstrate a fundamental limitation that mathematical and computational methods should be aware of in order to properly reconstruct the original multidimensional space from a receiver. To demonstrate this limitation, take for example received signals that encode (either in $1$D or $2$D signal streams) adjacency matrices of the (monoplex) networks whose isomorphic counterparts are multidimensional. For any arbitrarily chosen formal theory and decoding algorithm, the theorems in [35, 36] imply that there is a sufficiently large multidimensional network whose multidimensional space cannot be retrieved from the adjacency matrix of the network by the respective formal theory and/or decoding algorithm. This occurs because the algorithmic information distortions (due to being exponential) simply outgrow any prediction or output capability of the formal theory together with the decoding algorithm, should the network be given as input [36, Theorem 12].
@@
Such distortions demonstrate that in the general case where the multidimensional spaces may be much more complex than the objects themselves, there is a fundamental limitation for any formal theory with regard to the original multidimensional space to which a received message containing an object embedded into a very simple multidimensional space (e.g., a linear one, such as a signal stream) in fact corresponds [42]. This implies that even in the ideal scenario in which we correctly infer the actual object sent by a completely unknown emitter, we are still prone to being wrong about the original dimensions’ configurations the emitter intended. In addition to the distinctive feature of AID in comparison to classical AIT that we will discuss in Section A.5.1, these algorithmic distortions also pose a challenge to straightforward applications of the algorithmic coding theorem [38, 39, 40, 41], universal (Solomonoff) induction, or minimum description length methods [27, 40]. This challenge may be even more aggravated in the case of arbitrarily large signal streams which correspond to partial messages made available to the receiver agent so that the respective intended messages will be only completed in larger signal streams [43].
Intuitively, it is straightforward for us humans to conceive and construct an object, and then to transmit a signal with this object encoded in the message the signal is conveying, so that this object is much more complex than the space-time $4$-dimensional space that we are used to. However, in the general case either the multidimensional space of a unknown source can be much more complex than humans are used to or the object that the unknown emitter is trying to convey can be relatively simple in comparison to this yet unknown multidimensional space. Both cases can be affected by the limitations brought about by the aforementioned algorithmic information distortions.

### v3

@@ III-B Theoretical results
In the general case (where the multidimensional spaces may be much more complex than the objects themselves), even in the ideal scenario in which we correctly infer the actual object sent by a completely unknown emitter, we are still prone to being wrong about the original dimensions’ configurations the emitter intended [34, 35]. (See also Appendix A ) This poses a challenge to straightforward applications of the algorithmic information theory (AIT). In order to tackle such a challenge, we develop a method for multidimensional reconstruction of received low-dimensional signals in which one does not have access to knowledge about the encoding-decoding scheme or the original multidimensional space. Theorem 1 demonstrates the existence of sufficient conditions (or assumptions) that allow one to achieve an encoding/decoding-agnostic method for reconstructing the original multidimensional spaces. It gives rise to an agnostic method that does not need to assume that one a priori knows that the emitter agent’s multidimensional space resembles ours, is as complex as ours, or is isomorphic to the structures we currently deem fundamental in mathematics.
Theorem 1 (Theorem $2.10$ in the Sup. Mat.)
@@
hold, where $\mathcal{P}$ is any algorithmic perturbation that transforms $y$ embedded into $\mathcal{S}$ into another $y^{\prime}$ embedded into $\mathcal{S}^{\prime}$. Let $\mathbf{P}^{\prime}\left[\cdot,\cdot\right]$ be the uniform probability measure over the space of all possible outcomes (in the form $\left(y^{\prime},\,\mathcal{S}^{\prime}\right)$) of algorithmic perturbations on $\left(y,\,\mathcal{S}\right)$ that satisfy Equation 3. Let $\mathbf{P}\left[\cdot,\cdot\right]$ be any arbitrarily chosen (computable) probability measure on the space of all possible (computably constructible) objects embedded into a multidimensional space. Then, for any $\left(y^{\prime},\mathcal{S}^{\prime}\right)$ distinct from $\left(y,\mathcal{S}\right)$ satisfying Equation 3, the following strong asymptotic dominance
@@
holds with probability (given by the uniform probability measure $\mathbf{P}^{\prime}$) arbitrarily close to $1$ as $\left|y\right|\to\infty$.
@@
As further discussed in Appendices D and E, we believe the results in this article not only contribute with a general method for interpreting messages in zero-knowledge one-way communication channels, but also establish fundamental connections (and posit open problems) with zero-knowledge proofs and the algorithmic counterparts of the maximum likelihood method and the noisy-channel coding theorem.
III-C Toward a general theory for zero-knowledge one-way communication
The results presented in this article suggest that AID is capable of harnessing (given the availability of adequate computational resources) some of the algorithmic-informational properties that we demonstrated in ideal and generalised scenarios. As shown in [24, 40], AID allows the investigation of causal analysis [28, 29] and solution of inverse problems [25] by taking advantage of a high convergence rate to the algorithmic probability brought along by BDM, probability values which remain stable across distinct models of computation, most prominently for high algorithmically probable objects.
Our theoretical results show that the method employed to achieve the empirical findings in Section II-A—see also Sup. Mat. —is sound with regard to the infinite asymptotic limit when computational resources are unbounded; and it also shows that the method can be generalised to multidimensional spaces in addition to bi- or tridimensional ones. Our inference method is general enough to remain robust vis-a-vis any arbitrarily chosen alternative method to calculate the probabilities of the events, any arbitrarily chosen encoding-decoding scheme, programming language, or computation model, without the need of assuming some bias in the possible outcomes and prior beliefs.
Theorem 1 demonstrates that there are conditions for which decoding the message in the original multidimensional space $\mathcal{S}$ chosen from the one in which the message has the lowest algorithmic complexity (i.e., highest algorithmic probability, as assured by the universal distribution) becomes eventually more likely to be correct than decoding the message in a distinct multidimensional space $\mathcal{S}^{\prime}$, with or without noise. The underlying idea is that the larger the algorithmic complexity of the closer-to-randomness message candidate in the wrong multidimensional space with respect to the algorithmic complexity of the message in the correct multidimensional space, the larger the (algorithmic) likelihood [41, 42] of the correct multidimensional space with respect to the likelihood of the wrong closer-to-randomness multidimensional space (see also Appendix E). Thus, finding the multidimensional space that minimises the algorithmic complexity the most, eventually increases the probability of finding the correct one in comparison to the probability of finding the wrong closer-to-randomness one. Notice that this holds even if one assumes an arbitrarily different probability measure $\mathbf{P}$ as starting point.
In this manner, we demonstrate that once sufficient conditions for the compressibility of objects, multidimensional spaces, and perturbations (e.g., noise) are met, one-way communication to the receiver becomes possible even if no prior knowledge about the encoding-decoding scheme chosen by the emitter is known. One can first simply look into the landscape of all possible partitions and low-complexity noise the receiver has access to while assuming no bias toward particular objects and partitions in this landscape. Then, despite having no knowledge about the encoding-decoding schemes and thus a lack of biases, one can infer with probability as high as one wishes that the message with the lowest algorithmic complexity is arbitrarily more likely to be correct than any other arbitrary computable method for assigning probabilities to message candidates.
One of the main conditions for this to hold is that the algorithmic complexity of the object $y$ (with or without noise) in the wrong multidimensional space should increase (i.e., move the message toward algorithmic randomness) sufficiently fast in comparison to the algorithmic complexity of the message in a correct multidimensional space. To this end, one can assume a message should contain a large amount of redundancies, and thus should be highly compressible with respect to its linear signal stream length $\left|y\right|$, such as the case in which $\mathbf{K}\left(y\middle|\mathcal{S}\right)=\mathbf{O}\left(\log\log\left(\left|y\right|\right)\right)$ holds, where $\mathbf{K}\left(\cdot\right)$ denotes the algorithmic complexity. In this case, most low-complexity perturbations will transform the object and/or partition by sufficiently moving it toward algorithmic randomness so that the conditions are satisfied. This condition resembles the ones for the noisy-channel coding theorem in classical information theory [20], although our method can be understood as being more general than the classical version because: it holds for zero-knowledge communication channels, while the classical version requires that the encoding-decoding schemes must be previously known by the emitter and receiver; and stochastic noise (as in the classical version) is a particular case of randomly generated algorithmic perturbations [26]. See also Appendix C for further discussion on this and other mathematical phenomena.
The proofs of all theorems can be found in Section $2$ of the Sup. Mat..
IV Discussion
Our method is based on algorithmic information dynamics (AID) which combines tools from classical information theory and algorithmic information theory, in which algorithmic complexity values are approximated and combined with Shannon entropy by using the block decomposition method (BDM), in order to perform a perturbation analysis so that this algorithmic-information-based method can be used to avoid encoding multidimensional distortions or biases in signal detection and processing. Our method is asymptotically agnostic vis-a-vis not only encoding-decoding schemes, but also to the arbitrarily chosen: computation model or computational appliances made available; programming language; computable (or semi-computable) method of approximation to the algorithmic information content; (computable) probability measure of the events or models; and formal theory able to formalise the mathematical statements. Our theoretical framework and results can be extended and used for bio- and technosignature detection and signal deconvolution. Such results relate information theory to fundamental areas of mathematics, such as geometry and topology, by means of compression and algorithmic probability.
As shown in [24, 40], AID allows the investigation of causal analysis [28, 29] and solution of inverse problems [25] by taking advantage of a high convergence rate to the algorithmic probability brought along by BDM. In addition, although it may seem paradoxical at first glance—because one may initially think that the errors produced by the methods for approximating the algorithmic complexity would be inherited by the perturbation analysis phase itself—our theoretical and empirical results show that the perturbation analysis enables one to overcome the semi-computability limitation in algorithmic information theory, but only when, looking at the whole landscape of complexity values, one estimates with a sufficiently large amount of perturbations effected on the received signal. This is because the algorithmic complexities of the perturbed signals eventually become sufficiently larger than the complexity of the original message. Hence, a powerful enough method for approximating complexity, such as BDM, can eventually harness this divergence in order to enable the perturbation analysis to statistically distinguish the correct one (i.e., more compressible) from the wrong ones (i.e., more random). That is, an approximating method with enough computational resources can produce smaller errors than the difference of actual (although semi-computable) complexity values between the correct candidate and the perturbed ones, even though the exact complexity values remain ultimately unknown to the perturbation analyst (i.e., even though the analyst itself can never be sure it has enough computational resources in the first place). We argue this is related to zero-knowledge proofs and algorithmic generalised versions of the noisy channel coding and the maximum likelihood principle. Further discussion of these and other topics can be found in Appendices AE.
Acknowledgments
@@
Appendix A Avoiding algorithmic information distortions in arbitrarily complex multidimensional spaces
@@
In [34, 35] it is investigated how much of the irreducible information content of a multidimensional network and its isomorphic monoplex network (i.e., its isomorphic graph) is preserved during such an isomorphism transformation. In fact, isomorphisms are demonstrated to not preserve algorithmic information in the general case. The algorithmic information of a multidimensional network might be exponentially distorted with respect to the algorithmic information of its isomorphic (monoplex) network (see [35, Theorem 12]). We refer to an algorithmic information distortion when the algorithmic information content of an object (e.g., measured by the prefix algorithmic complexity of an encoded form of the object) in a certain context (e.g., the multidimensional space the object is embedded into) sufficiently differs from the algorithmic information content of the same object in a distinct context. In this regard, algorithmic complexity oscillations [38] may arise from a series of distortions produced by a large enough collection of distinct contexts into which the same object is embedded.
@@
As shown in [30] (see also Section $3$ in the Sup. Mat.) other examples are the oscillations measured by the plots presented in the figures, particularly those image plots for which one has the $x$-axis representing the distinct partitions we estimated the algorithmic complexity of the messages into, respectively. Notice that the downward spikes one can see in the figures are examples of algorithmic complexity (negative) oscillations such that for the decoding problem we are interested in, they indicate the partition that corresponds to the correct multidimensional space the message intended.
@@
The results in [34, 35] demonstrate a fundamental limitation that mathematical and computational methods should be aware of in order to properly reconstruct the original multidimensional space from a receiver. To demonstrate this limitation, take for example received signals that encode (either in $1$D or $2$D signal streams) adjacency matrices of the (monoplex) networks whose isomorphic counterparts are multidimensional. For any arbitrarily chosen formal theory and decoding algorithm, the theorems in [34, 35] imply that there is a sufficiently large multidimensional network whose multidimensional space cannot be retrieved from the adjacency matrix of the network by the respective formal theory and/or decoding algorithm. This occurs because the algorithmic information distortions (due to being exponential) simply outgrow any prediction or output capability of the formal theory together with the decoding algorithm, should the network be given as input [35, Theorem 12].
@@
Such distortions demonstrate that in the general case where the multidimensional spaces may be much more complex than the objects themselves, there is a fundamental limitation for any formal theory with regard to the original multidimensional space to which a received message containing an object embedded into a very simple multidimensional space (e.g., a linear one, such as a signal stream) in fact corresponds [43]. This implies that even in the ideal scenario in which we correctly infer the actual object sent by a completely unknown emitter, we are still prone to being wrong about the original dimensions’ configurations the emitter intended. Thus, in addition to the distinctive feature of AID in comparison to classical AIT that we will discuss in Appendix B, these algorithmic distortions also pose a challenge to straightforward applications of the algorithmic coding theorem [36, 37, 38, 39], universal (Solomonoff) induction, or minimum description length methods [27, 38]. If the complexity of the object itself is sufficiently smaller than the complexity of the space (or structure) the object is embedded into, there are cases in which minimising the overall algorithmic complexity (or maximising the algorithmic probability) of the entire message may lead to correctly guessing the object while incorrectly guessing the exact configuration of the (multidimensional) space. Such a challenge may be even more aggravated in the case of arbitrarily large signal streams which correspond to partial messages made available to the receiver agent so that the respective intended messages will be only completed in larger signal streams [44]. Intuitively, it is straightforward for us humans to conceive and construct an object, and then to transmit a signal with this object encoded in the message the signal is conveying, so that this object is much more complex than the space-time $4$-dimensional space that we are used to. However, in the general case either the multidimensional space of a unknown source can be much more complex than humans are used to or the object that the unknown emitter is trying to convey can be relatively simple in comparison to this yet unknown multidimensional space. Both cases can be affected by the limitations brought about by the aforementioned algorithmic information distortions.